# LED: Light Enhanced Depth Estimation at Night

### v2

@@ 3.2 Architectures
3.3 Learning Objectives
To train the encoder-decoder, we adopt a combination of loss functions inspired by [19]. The primary loss, $\mathcal{L}_{depth}$, is the Log L1 loss presented in Eq. 1. It measures the error between the estimated depth, $d_{i}$, and the corresponding ground truth, $g_{i}$. This variant of L1 loss attributes less significance to errors occurring at greater distances. This adjustment aligns with the expectation that given errors, e.g. 1 m, should have greater impact when within a few meters of the camera but are more tolerable at extended distances.
$\mathcal{L}_{depth}=\frac{1}{N}\sum_{i=1}^{N}|log(d_{i})-log(g_{i})|$ (1)
We address edges fidelity, particularly challenging in low-light conditions. To this end, we incorporate a loss, $\mathcal{L}_{grad}$, that specifically emphasizes on gradients. As we believe edges sharpness is important regardless of the distance, we employ a standard L1 Loss expressed in Sec. 3.3 instead of a logarithmic one used in [19]. Moreover, our experiments revealed no performance improvements with the Log version. $\nabla_{x}(d_{i})$ and $\nabla_{y}(d_{i})$ respectively represent the spatial derivative of $d_{i}$ along the x and y-axis.
$\mathcal{L}_{grad}=\frac{1}{N}\sum_{i=1}^{N}$ $|\nabla_{x}(d_{i})-\nabla_{x}(g_{i})|$
$+$ $|\nabla_{y}(d_{i})-\nabla_{y}(g_{i})|$ (2)
Similar to [19], we want to ensure accurate surfaces representation in depth maps. The depth normals are estimated at each pixel using $n_{i}^{a}\equiv[-\nabla_{x}(a_{i}),-\nabla_{y}(a_{i}),1]^{\intercal}$. The cosine similarity loss, expressed in Eq. 3, is then employed to compare estimated and ground truth normals. .,$\langle$.,$\rangle$ . , . ⟩ denotes vector inner product operation.
$\mathcal{L}_{normal}=\frac{1}{N}\sum_{i=1}^{N}|1-\frac{\langle n_{i}^{d},n_{i}^{g}\rangle}{\sqrt{n_{i}^{d},n_{i}^{d}}\sqrt{n_{i}^{g},n_{i}^{g}}}|$ (3)
Finally, our learning objective can be expressed as Eq. 4, we set $\lambda_{1}=1$ and $\lambda_{2}=1$.
$\mathcal{L}=\mathcal{L}_{depth}+\lambda_{1}\mathcal{L}_{grad}+\lambda_{2}\mathcal{L}_{normal}$ (4)
3.4 Implementation Details
We implement the encoder-decoder in PyTorch [31]. For each experiment the model is trained from scratch during 70 epochs using the AdamW optimizer [25]. The training is conducted with a batch size of 32, and both input and output resolutions are set to 320$\times$320 px. The learning rate is set to $10^{-3}$ and the selection of the best epoch is based on the Root Mean Square Error (RMSE) metric. This determination is made on a 5,000 images validation set. The entire training process takes approximately 4 hours, conducted on a single Nvidia RTX 4090.
Adabins was implemented using the official code from [4], while we rely on the Monocular-Depth-Estimation-Toolbox [22] for DepthFormer [23]. The only modifications are related to dataset format compatibility. The training setup aligns with the original papers’ provided recommendations, only excluding data augmentation to preserve the HD pattern from cropping.
Figure 4: Simulation of HD pattern: The top-left image corresponds to the control matrix of the HD headlight (black=0%, white=100% intensity). At the bottom-left, the generated photometry considering aberrations created by the headlight lens is displayed, particularly noticeable at the edges. On the right, an example of a simulated image using the photometry is presented. The area outlined in red corresponds to the region of interest.
Model RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
Encoder-decoder (HB) 6.1204 0.0903 0.0233 0.1353 13.4847 3.3579 0.9489 0.9812 0.9908
Encoder-decoder (LED) 5.4259 0.1996 0.0188 0.1253 12.4900 16.1224 0.9603 0.9846 0.9927
Adabins (HB) 7.3520 0.1360 0.0320 0.1230 10.1880 4.2770 0.8790 0.9440 0.9740
Adabins (LED) 5.5830 0.0690 0.0240 0.0920 7.7290 2.5090 0.9210 0.9680 0.9840
DepthFormer (HB) 5.8528 0.0687 0.0281 0.1205 11.3841 0.8328 0.9432 0.9802 0.9913
DepthFormer (LED) 5.3845 0.1215 0.0276 0.1111 10.4146 3.4015 0.9497 0.9838 0.9950
Table 1: Comparison of depth estimation performances: (HB) models are trained on High Beam data and (LED) on HD light pattern. (LED) models outperform (HB) models across all metrics, exceptions are for Abs Rel and Sq Rel. Metrics are computed in the ROI.
@@ 4 Dataset
4.1 Nighttime Synthetic Drive Dataset
Given the absence of public datasets containing HD pattern illumination, we create the Nighttime Synthetic Drive Dataset using Nvidia Drive Sim (Drop 15) simulator [30]. This simulator produces road images with photorealism effort. We adapt simulated vehicles headlights to enable real HD projections based on photometric measurements. Our generated dataset comprises 24,995 images with patterns and 24,995 images with high beam illumination. Examples are displayed in Fig. 3.
We select high beam (HB) as comparison point, considering it to be the maximum normal lighting condition. Moreover, high beam provides more illumination than the checkerboard pattern, ensuring that performance improvement are only due to the pattern contribution. All images are simulated nighttime scenes.
Elements in the scene, such as vehicles, pedestrians, and traffic signs, are randomized in each frame. To account for possible light interference, we simulate other car headlights and randomize ambient light between 0 and 10 lux.
We use five different maps and generate 4,999 frames for both pattern and HB illuminated images in each. Maps are intended to represent major cities in specific regions of the world. Three of them are used for training, one for validation, and one for testing. While the two parts of the dataset are not strictly identical due to randomization of road entities, they share the same maps and randomization code, making the domains comparable.
To approach reality as closely as possible, we account for headlight’s optical imperfections using real measured photometry file. Fig. 4 illustrates differences between the ideal control matrix and real projection applied in simulation. Thus, aberrations observed in the pattern, particularly along the edges of the projection, are indicative of these optical imperfections.
All RGB images have a 1920$\times$1080 px resolution and corresponding depth maps. Furthermore, we provide annotations necessary for object detection 2D/3D, normals estimation, semantic and instance segmentation, making the dataset suitable for a multitude of tasks related to autonomous driving.
As we are not constrained by sensor capabilities, the provided data are not limited, yet we clip the depth at 100 m for all experiments. For reference, KITTI [14] and Oxford RobotCar [27], two real-world road datasets widely used, are respectively limited to 90 m and 50 m. During training, a 640$\times$640 px square is initially center-cropped and then resized to 320$\times$320 px.
To ensure results reproducibility, code and synthetic dataset are publicly available. Releasing large-scale nighttime data, with precise and comprehensive annotations, will foster research on nighttime computer vision for autonomous driving.
@@ 4.2 Real-world dataset
We collected an in-house dataset using a real car-mounted prototype to demonstrate the applicability of our method to real-world scenario. It comprises 50,000 images (70% train / 15% val / 15% test) from populated urban and rural roads, evenly split between low beam (LB) illumination and checkerboard pattern with 0.5°, 0.25° and 0.125° cells’ dimension. It has been annotated from aggregated and densified LiDAR data using Exwayz software [12]. The HD headlamp we used offers greater projection resolution than the simulated one. More details are available in supplementary materials.
@@ 5 Experiments
We evaluate our method through extensive experiments on the Nighttime Synthetic Drive Dataset. We demonstrate contribution of our light pattern in boosting depth estimation performance, both inside and outside illuminated area. We apply our method to other SOTA approaches: Adabins [4] and DepthFormer [23], showing its extensibility. Finally, we evaluate LED robustness beyond its training domain and its capabilities on real-world scenarios.
The metrics we use in our experiments align with previous works [4, 23, 38, 24, 45, 47]: Root Mean Squared Error (RMSE), Absolute Relative error (Abs Rel), Logarithmic error ($\text{Log}_{10}$), RMSE in Log Space (RMSE Log), Scale-Invariant Logarithmic error (SILog), Squared Relative error (Sq Rel), and threshold accuracy ($\delta^{1}$, $\delta^{2}$, and $\delta^{3}$) are employed for a comprehensive evaluation of depth estimation performance.
Figure 5: Qualitative results. LED results (right) exhibit higher precision, showcasing more accurate object boundaries and shapes compared to HB (middle). Red boxes indicate enhanced regions.
Model RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
@@
Table 2: Encoder-decoder performance beyond ROI: O-ROI stands for Outside ROI, where the evaluation mask is the inverse of the ROI. For Full, metrics calculation is performed on the entire image. (LED) models trained on HD pattern consistently outperform (HB) models in all metrics.
@@ 5.1 HD Pattern Impact
We showcase the depth estimation enhancement obtained by comparing results with and without pattern. As reference we take the encoder-decoder (Sec. 3.1), trained on high beam (HB), while the variant uses our HD pattern (LED). Quantitative and qualitative results are respectively presented in Tab. 1 and Fig. 5.
Quantitative Results.
Comparing LED with HB (Tab. 1), we observe a significant improvement in metrics within the ROI: LED yields -11% RMSE. The introduction of the checkerboard pattern leads to a degradation in relative metrics (Abs Rel, Sq Rel). Indeed, when using LED, the model provides a better estimation of distances. However, errors are often related to pattern occlusions, indiscriminately made on both near and far objects, which impact relative metrics. All metrics assessing the global precision show improvements.
In comparison with the two SOTA approaches [4, 23] mentioned in Tab. 1, LED-trained encoder-decoder either outperforms or matches the performance of their models trained on HB. Considering that the encoder-decoder architecture is inherently less tailored for depth estimation compared to SOTA more intricate designs, this result underscores the substantial improvements yields by LED. Incorporating HD informative pattern, even with straightforward architecture, is therefore promising for challenging nighttime scenarios.
Performance over Distance.
Due to the limited visibility at nighttime, depth estimation of distant objects is challenging. In Fig. 6 (a), we showcase the performance of LED over varying distances compared to HB. While maintaining similar performance for close objects ($<$20m), LED exhibits greater enhancement at longer ranges. Thus, it effectively addresses the task of depth estimation for distant objects at night.
@@
Table 3: Performances using various pattern on half-sized dataset. VL and HL refers to vertical and horizontal lines
Pattern Study.
We generated additional data with various illumination pattern to compare performances between LED checkerboard and other common structured light patterns (horizontal and vertical lines). Tab. 3 demonstrates that choosing the checkerboard pattern leads to the best performance across most metrics. Horizontal lines exhibit slight improvement in relative metrics but a significant decline in other metrics.
Domain RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
@@
Table 4: Encoder-decoder performance across domains: We denote (training domain $\rightarrow$ testing domain). The model trained on (HB) fails entirely when tested on (LED). Yet, (LED)-trained models have the ability to estimate depth when tested on the (HB) domain.
ROI-only Training.
We investigate whether training the network exclusively on the ROI could potentially enhance our method. This strategy results in an additional improvement of -8.58% Abs Rel, offset by +4.50% RMSE within the specified region, compared to the full-image trained network. This trade-off is valuable for applications where Abs Rel holds greater significance. Additional results can be found in the supplementary material.
Figure 6: We compare LED to HB in various scenarios. (a) Performance over distance; (b) Performance over training set size; (c) Robustness across domains: Number of training data is fixed and composed of a given ratio of LED and HB data. Overall, LED achieves better long-range results while being more data-efficient. It also demonstrates improved performance across domains.
@@ 5.2 Improving Global Scene Understanding
We investigate the impact of LED on depth estimation beyond ROI. Observed improvement is an indication of the pattern’s contribution to a more holistic scene understanding. Results are presented in Tab. 2.
Performance outside the ROI.
Examining depth estimation only outside the ROI, we observe -2.66% improvement in RMSE. We note the absence of degradation in relative metrics, unlike results within the ROI presented in Sec. 5.1. Instead, we report -3.03% Abs Rel and -23.27% Sq Rel. This suggests that the HD headlight pattern provides valuable information for a more accurate understanding of the broader scene geometry. This includes insights into object sizes and scale.
Performance outside the entire image.
Expanding our analysis to the entire image - ROI and outside ROI - we witness improvements such as -3.05% RMSE. Similar to outside ROI, we note -1.50% Abs Rel and -9.23% Sq Rel. This indicates that the HD headlight pattern contributes significantly to the global scene understanding, offering valuable cues that benefit overall depth estimation.
@@ 5.3 Annotation-efficient learning
Collecting annotated nighttime data is both challenging and costly, thus, we investigate whether extra guidance provided by the pattern could lead to reduced needs of training data (refer to Fig. 6 (b)). We demonstrate that encoder-decoder trained with less than 20% of the LED data outperforms model trained on the entire HB dataset. Furthermore, model’s performance is near its peak even when trained with only 50% of the dataset.
In addition, we observe that incorporating just 10% of LED data into the HB training set enable the network to learn relevant features and enhance its performances when used with pattern (see Fig. 6 (c)). Therefore, any vehicle equipped with HD headlights can apply LED by adding only a few pattern images in their training set, thereby reducing the cost of specialized data.
Model RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
Encoder-decoder (LB) 14.199 0.187 0.072 0.292 28.938 5.620 0.772 0.909 0.956
Encoder-decoder (LED 0.5°) 11.089 0.119 0.050 0.213 21.251 2.304 0.864 0.950 0.977
Encoder-decoder (LED 0.25°) 8.695 0.109 0.040 0.190 18.635 2.128 0.899 0.963 0.984
Encoder-decoder (LED 0.125°) 10.154 0.096 0.040 0.190 18.862 2.132 0.907 0.969 0.984
DepthFormer (LB) 8.777 0.138 0.050 0.186 17.170 3.365 0.864 0.957 0.983
DepthFormer (LED 0.5°) 6.810 0.101 0.041 0.152 14.380 1.376 0.890 0.969 0.989
DepthFormer (LED 0.25°) 5.621 0.082 0.030 0.126 11.735 1.163 0.923 0.978 0.992
DepthFormer (LED 0.125°) 5.727 0.076 0.030 0.116 10.754 1.261 0.940 0.987 0.995
Table 5: Performance comparison on our real-world dataset using various pattern resolutions: (LED) models outperform (LB) across all metrics, demonstrating LED applicability on real-world scenarios.
Figure 7: Real world qualitative results: DepthFormer LED model produces accurate depth estimation on complex scenes, even in presence of interfering lights, e.g., car headlights, street lights.
5.4 Extensibility to Other SOTA Methods
We extend the evaluation of LED beyond the encoder-decoder architecture, by replicating the methodology of Sec. 5.1 with two additional SOTA architectures. Quantitative results are available in Tab. 1.
As we provide a new annotated dataset, we opt for two high performing SOTA supervised approaches designed for daytime scenarios: Adabins [4] and DepthFormer [23]. Their use of transformer architectures provides valuable insights into our method behavior when used on more complex architectures.
With the HD pattern, Adabins demonstrates a substantial improvement with -24.06% RMSE and -6.70% Abs Rel within the ROI compared to the HB model. Conversely, DepthFormer showcases -8.00% enhancement in RMSE but +5.28% decrease in Abs Rel. Note that the most significant enhancement is observed in Adabins, which initially had the worst results by night with HB.
Improvements seen across diverse architectures confirm LED agnosticity to architectures. It implies potential effectiveness with future SOTA methods.

### v3

@@ 4 Dataset
4.1 Nighttime Synthetic Drive Dataset (NSDD)
[…]
We choose High Beam (HB) as the comparison baseline, considering it the maximum normal lighting condition. Since high beam illumination exceeds that of the checkerboard pattern, performance improvements are attributed solely to the pattern’s contribution. In each frame, vehicles, pedestrians, and traffic signs are randomized. To simulate light interference, we include other car headlights and randomize ambient light levels between 0 and 10 lux.
Figure 4: Simulation of HD pattern: Top-left: control matrix of the HD headlight. Bottom-left: photometry considering aberrations created by the headlight lens. Right: Resulting image using the photometry. The area outlined in red is the region of interest.
The dataset consists of 5 different maps (3 train / 1 val / 1 test), with 4,999 frames generated for both pattern and HB illuminated images in each. These maps represent major cities. Both parts of the dataset share the same randomization code, making the domains comparable.
[…]
@@ 4.2 Real-world Dataset
We validate LED on a real-world, in-house dataset collected using a car-mounted prototype. It comprises 50,000 images (70% train / 15% val / 15% test) from populated urban and rural roads, evenly split between Low Beam (LB) illumination and checkerboard pattern with 0.5°, 0.25° and 0.125° cells’ dimension. Ground truth is obtained from LiDAR data using the DOC-Depth method [de Moreau et al.(2025)de Moreau, Mathias, Hassan, Yasser, Andrei, Hafid, Fabien, et al.] and Exwayz software [Exwayz(2024)]. More details are available in supplementary materials.
@@ 5 Experiments
Figure 5: Qualitative results on NSDD. LED results exhibit higher precision, and more accurate object boundaries and shapes compared to HB. Red boxes indicate enhanced regions.
Pattern RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
[…]
HB 6.1204 0.0903 0.0233 0.1353 13.4847 3.3579 0.9489 0.9812 0.9908
LED 5.4259 0.1996 0.0188 0.1253 12.4900 16.1224 0.9603 0.9846 0.9927
[…]
HB 7.3520 0.1360 0.0320 0.1230 10.1880 4.2770 0.8790 0.9440 0.9740
LED 5.5830 0.0690 0.0240 0.0920 7.7290 2.5090 0.9210 0.9680 0.9840
[…]
HB 5.8528 0.0687 0.0281 0.1205 11.3841 0.8328 0.9432 0.9802 0.9913
LED 5.3845 0.1215 0.0276 0.1111 10.4146 3.4015 0.9497 0.9838 0.9950
Table 1: Comparison of depth estimation performances on NSDD: HB models are trained on high beam data and LED on HD light pattern. LED models outperform HB models across all metrics, exceptions are for Abs Rel and Sq Rel. Metrics are computed in the ROI.
We evaluate our method through extensive experiments on the Nighttime Synthetic Drive Dataset. We validate the contribution of our light pattern in boosting performance, both inside and outside illuminated area. We apply our method to the encoder-decoder and other SOTA approaches: Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka], DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang] and Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], showing its agnosticity. Finally, we evaluate LED robustness beyond its training domain and its capabilities on real-world scenarios. Metrics used align with prior works [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Li et al.(2023)Li, Chen, Liu, and Jiang, Spencer et al.(2020)Spencer, Bowden, and Hadfield, Liu et al.(2021)Liu, Song, Wang, Liu, and Zhang, Wang et al.(2021)Wang, Zhang, Yan, Li, Xu, Li, and Yang, Zheng et al.(2023)Zheng, Zhong, Li, Gao, Zheng, Jin, Wang, Zhao, Zhou, Zhang, et al.].
Model RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
@@
Table 2: Encoder-decoder performance beyond ROI on NSDD: O-ROI stands for Outside ROI, where the evaluation mask is the inverse of the ROI. For Full, metrics are computed on the entire image. (LED) models trained on HD pattern outperform (HB) ones in all metrics.
Figure 6: Performance comparison of LED and HB in various settings and metrics on NSDD. (a) Performance over distance; (b) Performance over training set size; (c) Robustness across domains: The amount of training data is fixed and composed of a given ratio of LED and HB data. Overall, LED achieves better long-range results while being more data-efficient. It also demonstrates improved performance across domains.
@@ 5.1 HD Pattern Impact
We assess LED impact on depth estimation by comparing results with and without pattern. As reference, we take models trained on high beam (HB), while the variants use our HD pattern (LED). Quantitative and qualitative results are presented in table˜1 and fig.˜5.
Quantitative Results. We evaluate our method on the encoder-decoder and two SOTA architectures commonly used in daytime scenarios, Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka] and DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang]. Their use of transformer architectures offers valuable insights into LED behavior on more recent architectures compared to the convolutional ones that are common in embedded devices.
Comparing the encoder-decoder LED with HB (table˜1), we observe a significant improvement in metrics within the region-of-interest (ROI): LED yields -11% RMSE. Using LED, the model provides more accurate distance estimations. However, its errors are often related to pattern occlusions, which affect both near and far objects, resulting in a degradation of relative metrics (Abs Rel, Sq Rel). All metrics assessing global precision are improved.
Compared to the SOTA approaches [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Li et al.(2023)Li, Chen, Liu, and Jiang], LED-trained encoder-decoder either outperforms or matches the performance of their HB models. Given that the encoder-decoder is less tailored for depth estimation than more intricate SOTA designs, this result underscores the substantial improvements yield by LED, demonstrating that using the HD informative pattern even with a straightforward architecture is promising for challenging nighttime scenes.
With LED, Adabins demonstrates a substantial improvement compared to the HB model: -24.06% RMSE and -6.70% Abs Rel. Conversely, DepthFormer showcases -8.00% in RMSE but +5.28% in Abs Rel. Note that the most significant enhancement is observed in Adabins, which have the worst results by night on HB.
Improvements seen across diverse architectures confirm LED agnosticity to architectures. It implies potential effectiveness with future methods.
Performance Over Distance. Limited nighttime visibility makes depth estimation of distant objects challenging. In fig.˜6 (a), we show the performance of the encoder-decoder over distances. LED maintains performance for close object and exhibits greater enhancement at longer ranges, thus addressing depth estimation of distant objects at night.
@@
Table 3: Performances of the encoder-decoder using various patterns on NSDD. VL and HL refers to vertical and horizontal lines. Results are obtained from 50% of NSDD training data.
Pattern Study. We generated additional data with various illumination patterns to compare the performance of the LED checkerboard against other common structured light patterns (horizontal and vertical lines). Table˜3 shows that the checkerboard pattern outperforms the others across most metrics. While horizontal lines show slight improvement in relative metrics, they cause a significant decline in others.
Domain RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
@@
Table 4: Encoder-decoder performance across domains of NSDD. We denote (training domain $\rightarrow$ testing domain). The model trained on (HB) fails entirely when tested on (LED). Yet, (LED)-trained models are able to estimate depth when tested on the (HB) domain.
Pattern # Images RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
HB Zero-shot 19.058 0.430 0.243 0.587 27.004 8.134 0.086 0.186 0.785
LED 18.023 0.420 0.232 0.563 25.500 7.779 0.096 0.214 0.867
HB 100 6.499 0.075 0.033 0.133 12.894 0.992 0.929 0.977 0.991
LED 6.198 0.073 0.032 0.131 12.182 0.898 0.933 0.977 0.991
HB 500 5.447 0.058 0.025 0.112 10.629 0.688 0.950 0.985 0.995
LED 5.225 0.057 0.025 0.108 10.038 0.648 0.953 0.986 0.995
HB 1000 5.124 0.058 0.026 0.107 10.038 0.607 0.953 0.985 0.995
LED 4.328 0.044 0.020 0.089 8.374 0.430 0.962 0.991 0.997
Table 5: Depth Anything V2 performances on NSDD. The model fails when used as zero-shot, but it correctly estimates depth with few-shot learning. LED always outperforms HB.
@@ 5.2 Improving Global Scene Understanding
We investigate the impact of LED on depth estimation beyond the ROI, highlighting its contribution to a more holistic scene understanding. Results are presented in table˜2.
Examining depth estimation both outside the ROI and across the entire image, we observe a consistent performance improvement: -2.66% RMSE outside the ROI and -3.05% RMSE on the entire image. We also report enhancement in relative metrics: Abs Rel (resp. -3.03% and -1.50%) and Sq Rel (resp. -23.27% and -9.23%). These results suggest that the HD headlight pattern provides valuable information, such as object size and scale, leading to enhanced overall scene understanding and more accurate depth estimation.
@@ 5.3 Annotation-efficient Learning
Collecting annotated nighttime data is challenging and costly. We explore whether pattern extra guidance can reduce the need for training data (refer to fig.˜6 (b)). Our results show that encoder-decoder model trained with less than 20% of the LED data outperforms one trained on the full HB dataset, reaching near-peak performance with just 50%. In addition, we observe that incorporating just 10% of LED data into the HB training set enable the network to learn relevant features and enhance its performances when used with pattern (see fig.˜6 (c)). Therefore, any vehicle equipped with HD headlights can apply LED by adding only a few pattern images in their training set, thereby reducing the cost of specialized data.
5.4 Few-shot Learning
Foundation models show strong zero-shot performance across various cameras and content types. However, when tested on NSDD, Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] performs poorly, with an RMSE of $\sim$19 m (see table˜5), highlighting the complexity of nighttime data and the need for specialized models. It performs slightly better on LED images, possibly due to its ability to leverage the pattern’s geometric cues. Fine-tuning with as little as 100 images is sufficient for adapting the network to our distribution and produce consistent depth predictions on both LED and HB. LED-models demonstrate significant improvements across all metrics (-15.5% RMSE with 1,000 images). This underlines the effectiveness of our method in enhancing nighttime depth estimation even on foundation models. Results on our real-world dataset are provided in the supplementary material.
Pattern RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
[…]
LB 14.199 0.187 0.072 0.292 28.938 5.620 0.772 0.909 0.956
LED 0.5° 11.089 0.119 0.050 0.213 21.251 2.304 0.864 0.950 0.977
LED 0.25° 8.695 0.109 0.040 0.190 18.635 2.128 0.899 0.963 0.984
LED 0.125° 10.154 0.096 0.040 0.190 18.862 2.132 0.907 0.969 0.984
[…]
LB 8.777 0.138 0.050 0.186 17.170 3.365 0.864 0.957 0.983
LED 0.5° 6.810 0.101 0.041 0.152 14.380 1.376 0.890 0.969 0.989
LED 0.25° 5.621 0.082 0.030 0.126 11.735 1.163 0.923 0.978 0.992
LED 0.125° 5.727 0.076 0.030 0.116 10.754 1.261 0.940 0.987 0.995
[…]
LB 7.637 0.104 0.044 0.160 14.617 1.401 0.891 0.972 0.990
LED 0.5° 7.017 0.109 0.047 0.161 14.044 1.260 0.877 0.9697 0.989
LED 0.25° 6.304 0.102 0.044 0.146 12.300 1.079 0.885 0.973 0.993
LED 0.125° 5.785 0.082 0.037 0.126 10.673 0.840 0.921 0.981 0.992
Table 6: Performance comparison on real-world data using various pattern resolutions. LED models outperform LB on all metrics, showing LED suitability on real-world scenarios.